OpenStudy (anonymous):

sqrt{x + 2} = x - 4 Find x (which is 7).

OpenStudy (anonymous):

@math0101 Square both sides, can you?

OpenStudy (anonymous):

x + 2 = x2 - 8x + 16

OpenStudy (anonymous):

Great :) Now simplify it into quadratic form >)

OpenStudy (anonymous):

hmm x + 2 = (x - 4 )^2

OpenStudy (anonymous):

still stuck :)

OpenStudy (anonymous):

x^2 - 8x - x + 16 -2 = 0 Add them together

OpenStudy (anonymous):

aha! x^2 - 9x -14 = 0 x = 9/2 +- sqrt{(9/2)^2 + 14} x = 9/2 0- 5/2 x1 = 7 x2 = 2 x1 is correct, but for some reason my book doesn't mention x2

OpenStudy (anonymous):

x = 9/2 0- 5/2 SHOULD READ: x = 9/2 +- 5/2

OpenStudy (anonymous):

I think I'll mark this as done. :) Thank you!

OpenStudy (anonymous):

Now pay attention to validate the solutions in square root case!

OpenStudy (anonymous):

oh yes... so much to remember! Like learning a language ;) Thanks!

OpenStudy (anonymous):

Domain: x ≥ 4